Right triangle DEF has sides that measure x, x, and x squared for the some value of x. What is the area of triangle DEF to the nearest 0.1 square unit?
Answer by richwmiller(17219)

x^2+x^2=x^4
2x^2=x^4
2=x^2
sqrt(2)=x